OpenStudy (gabylovesyou):

A study of three hundred teenagers found that the number of hours they spend on instant messaging sites each week is normally distributed with a mean of 13 hours. The population standard deviation is 5 hours. What is the margin of error for a 99% confidence interval? 0.743 0.334 0.320 0.413 @John_ES

OpenStudy (gabylovesyou):

B ?

OpenStudy (gabylovesyou):

@John_ES

OpenStudy (emilywandsnap132):

first why do u think it is B

OpenStudy (emilywandsnap132):

gtg

OpenStudy (gabylovesyou):

nvm i dont know how to do this cx so forget B ..

OpenStudy (john_es):

Do you know something about normal distribution?

OpenStudy (gabylovesyou):

no idea

OpenStudy (john_es):

Ok. Let's remember some concepts.

OpenStudy (john_es):

First the confidence interval for a normal distribution, \[(\overline{x}-Z_{\alpha/2}\frac{\sigma}{\sqrt{n}},\overline{x}+Z_{\alpha/2}\frac{\sigma}{\sqrt{n}})\]

OpenStudy (john_es):

Where, \[\overline{x}\] is the mean, \[Z_{\alpha/2}=2.575\]for a confidence level of 99%, \[\sigma\] is the standard deviation, and \[n\] is the size of the sample. Ok?

OpenStudy (john_es):

You need the margin error. I would say that the error can be obtained from the confidence interval, using this formula, \[Error=Z_{\alpha/2}\frac{\sigma}{\sqrt{n}}\]

OpenStudy (john_es):

As you see, is the value that is displaced from the mean.

OpenStudy (gabylovesyou):

ok...

OpenStudy (john_es):

Now, you can substitute, \[n=100; \sigma=5; \ Z_{\alpha/2}=2.575\]

OpenStudy (john_es):

Sorry, n=300

OpenStudy (john_es):

Then you have, \[E=2.575\cdot 5/\sqrt{300}=0.743\] The first answer.